# Empirical Best Prediction of Poverty Indicators via Nested Error Regression with High Dimensional Parameters

### v1

@@ Abstract
This paper extends the Nested Error Regression Model with High-Dimensional Parameters (NERHDP) to address challenges in small area poverty estimation. Building on the NERHDP framework, we develop a robust and flexible approach to derive empirical best predictors (EBPs) of small area poverty indicators, while accommodating heterogeneity in regression coefficients and sampling variances across areas. To overcome computational limitations in the existing algorithm, we introduce an efficient method that significantly reduces computation time, enhancing scalability for large datasets. Additionally, we propose a novel method for generating area-specific poverty estimates for out-of-sample areas, improving the reliability of synthetic estimates. To quantify uncertainty, we introduce a parametric bootstrap method tailored to the extended model. Through design-based simulation studies, we demonstrate that the proposed method has better performance in terms of relative bias and relative root mean squared prediction error compared to existing approaches. Furthermore, the proposed method is applied to household survey data from the 2002 Albania Living Standards Measurement Survey to estimate poverty indicators for 374 municipalities using auxiliary information from the 2001 census.
Poverty mapping,
small area estimation,
data linkage,
robust inference,
keywords:
[…]
, and
@@ 1 Introduction
The United Nations emphasizes the priority of eradicating all forms of poverty by making it the first of its 17 Sustainable Development Goals. Measuring poverty levels across regions is essential for designing effective poverty reduction strategies and ensuring equitable allocation of resources. Accurate information about people’s living conditions at a regional level is a basic instrument for targeting policies and programs to reduce poverty. However, progress toward this goal is often slowed by a consistent lack of data on important economic indicators, especially for small population groups and specific geographic areas in developing countries. One major reason for the shortage of of disaggregated poverty data is that large scale sample surveys are typically designed to produce reliable estimates of various characteristics of interest for large geographic areas or large population subgroups, rather than for smaller, more granular domains (Chatterjee, Lahiri and Li, 2008).
@@
Many measures of poverty and inequality are nonlinear functions of the small area distribution of the economic variable underpinning poverty, usually income or a surrogate such as expenditure. In this paper, we focus on estimating a class of poverty measures known as the Foster–Greer–Thorbecke (FGT) poverty measures, introduced by Foster, Greer and Thorbecke (1984). Consider a finite population of size $N$ partitioned into $m$ small areas of size $N_{1},\cdots,N_{m}$. Let $E_{ij}$ be a suitable quantitative measure of welfare for individual $j$ in small area $i$, such as income or expenditure, and let $z$ be a fixed poverty line. This poverty line $z$ serves as the threshold below which an individual is considered to be living in poverty. The FGT family of poverty measures for each small area $i$ is then defined as the area-specific mean
@@
The small area estimation (SAE) literature for poverty measures was revolutionized by the pioneering work of Elbers, Lanjouw and Lanjouw (2003), which introduced a unit-level approach (ELL) for poverty mapping. This method has become widely adopted for producing fine-grained estimates of FGT measures. The ELL approach assumes a two-level model with a common regression relationship between the response and auxiliary variables while allowing for heteroskedastic unit-level errors. The unit-level variance components are estimated using a proposed alpha model. Simulated census data are then generated to estimate poverty measures. It is important to note that the ELL approach is a synthetic method that does not rely on a conditional distribution. For further details, we refer readers to their original paper.
@@
Another popular approach for poverty estimation introduced by Molina and Rao (2010) is the empirical best prediction (EBP) method through Monte Carlo simulations. This method assumes that the transformed population welfare variables is characterized by a homoskedastic nested error regression (NER) model, as proposed by Battese, Harter and Fuller (1988). Specifically, it assumes that the regression coefficients and variance components are the same across all small areas, with normally distributed errors. Model parameters are estimated using the maximum likelihood (ML) method or the restricted maximum likelihood (REML) method, and EBPs of poverty indicators are then approximated through random draws from the conditional distribution of the out-of-sample units, given the sample data. Molina, Nandram and Rao (2014) introduced a Bayesian version of this approach, which is particularly useful for quantifying the uncertainty of the estimates.
@@
The effectiveness of such EBP method tends to decline when the number of small areas to be combined becomes large. This limitation arises because the assumption of identical regression coefficients and/or variance components across all small areas, inherent in these models, may not hold universally, leading to potential issues with model misspecification. To address this, random area-specific regression coefficient models (Prasad and Rao, 1990; Hobza and Morales, 2013; Rao and Molina, 2015) and random area-specific sampling variance models (Otto and Bell, 1995; Arora, Lahiri and Mukherjee, 1997; Kubokawa et al., 2016; Sugasawa and Kubokawa, 2017) have been suggested in the literature. Such modeling, though useful in some applications, needs more nontrivial assumptions on the joint distribution for a large number of random effects. On the other hand, fixed effects assumptions on the area specific regression coefficients and sampling variances generally lead to unstable estimates of these fixed effects due to small area specific sample sizes (Jiang and Nguyen, 2012).
@@
Lahiri and Salvati (2023) introduced a novel approach in the SAE literature, addressing several challenges. Specifically, this approach assumes fixed effects for both area-specific regression coefficients and sampling variances. But they estimate these coefficients using area-specific (with respect to a tuning parameter) estimating equations applied to data from all areas, and then use appropriately constructed residuals for estimation of variance components. When the tuning parameters are known, they haven shown that the model parameters can be consistently estimated. This approach, termed as the Nested Error Regression Model with High-Dimensional Parameters (NERHDP), not only offers a more flexible SAE modeling framework but also provides a reliable solution to the challenges of high-dimensional parameter estimation in small area analysis. However, their study focused solely on empirical best predictors (EBPs) for small area means, which are linear functions of the data.
@@
In this paper, we extend the NERHDP model to obtain EBPs of poverty indicators, as particular complex and nonlinear FGT measures. Unlike the approach in Molina and Rao (2010), our proposed EBPs account for heterogeneity across small areas in both regression coefficients and sampling variances. Such heterogeneity is common in real-world applications due to variations in socio-economic conditions, data quality, and sampling designs among areas. By extending NERHDP to accommodate FGT poverty measures, we enable the generation of more precise and reliable estimates that are sensitive to area-specific characteristics. This innovation not only enhances the flexibility and applicability of the modeling but also provides a robust methodological framework for addressing the complexities inherent in small area poverty estimation.
@@
Additionally, we address the computational challenges associated with the algorithm proposed by Lahiri and Salvati (2023), which involves iterative convergence and can be computationally intensive. We introduce a more efficient algorithm that significantly reduces computation time, producing estimates within seconds. This improvement ensures the proposed model is not only methodologically robust but also practical for large-scale applications.
@@
We conduct design-based simulation studies using the 2021 Italian dataset from the European Survey on Income and Living Conditions as the fixed population, which exhibits potential heterogeneity in both regression coefficients and sampling variances, to evaluate the performance of our proposed method. We compare it with existing approaches, including the EBP method of Molina and Rao (2010) (MR), the synthetic method of Elbers, Lanjouw and Lanjouw (2003) (ELL), and the direct estimator. The simulation results demonstrate that our method has better performance than the existing approaches in estimating headcount ratios and poverty gaps for both in-sample and out-of-sample areas, based on relative bias (RB) and relative root mean squared prediction error (RRMSPE).
We demonstrate the benefits of implementing the proposed approach on data from the 2002 Living Standards Measurement Survey (LSMS) in Albania to estimate headcount ratios and poverty gaps for the 374 municipalities, using auxiliary information from the 2001 Census. Of these municipalities, $161$ are out-of-sample, meaning they were not included in the original probability sample used for estimation. Consequently, their characteristics were not directly observed in the survey, and any estimates for these areas rely entirely on model-based predictions.
The structure of the paper is as follows. Section 2 provides a brief review of the empirical best prediction (EBP) methodology for FGT measures. Section 3 introduces the 2002 Living Standards Measurement Survey (LSMS) in Albania, describes the estimation problem, and outlines the available auxiliary information. Section 4 introduces the best prediction approach under a nested error linear regression model with high-dimensional parameters, along with the associated new parameter estimation algorithms. In Section 5, we describe the parametric bootstrap method used to estimate measures of uncertainty. Section 6 presents the results of design-based simulation studies, comparing the performance of the proposed method with existing approaches in terms of bias and variability. In Section 7, we apply the new methodology to LSMS data to estimate headcount ratios (HCRs) and poverty gaps (PGs) at the municipality level. Finally, Section 8 concludes the paper by summarizing the key findings and suggesting directions for future research.
@@ 2 A review of Empirical Best Prediction for FGT measures
Like in Molina and Rao (2010), we assume that a random sample of size $n<N$ is drawn from the finite population using a specified sampling design. Let $\Omega$, $s\subset\Omega$, and $r=\Omega-s$ denote the set of indices for all population units, the set of sampled units, and the set of indices for the unsampled units with size $N-n$, respectively. For area $i$, we have the following associated notations: $\Omega_{i}$, $s_{i}$, and $n_{i}$. Note that if area $i$ is not sampled, then $n_{i}=0$.
@@
where the expectation is taken with respect to the conditional distribution of $\mathbf{y_{r}}$ given the observed sample data $\mathbf{y}_{s}$.
@@
[…]
@@
Due to the complexity of $h_{\alpha}(t)$, it may not be feasible to explicitly calculate the expectation in (5). Assuming that $\mathbf{y}=(\mathbf{y}_{s}^{\prime},\mathbf{y}_{r}^{\prime})^{\prime}$ follows a multivariate normal distribution with mean vector $\bm{\mu}=(\bm{\mu}_{s}^{\prime},\bm{\mu}_{r}^{\prime})^{\prime}$ and a covariance matrix partitioned as:
@@
$\mathbf{y}_{r}|\mathbf{y}_{s}\sim N(\bm{\mu}_{r|s},\mathbf{V}_{r|s}),$ (6)
@@
Since the mean vector $\bm{\mu}$ and covariance matrix $\mathbf{V}$ typically depend on an unknown parameter vector $\bm{\phi}$, the conditional density $f_{Y_{ij}}(t|\mathbf{y}_{s})$ also depends on $\bm{\phi}$. Substituting $\bm{\phi}$ with its estimate $\hat{\bm{\phi}}$, we generate simulated values $Y_{ij}^{(k)}$ from the estimated density $f_{Y_{ij}}(t|\mathbf{y}_{s},\hat{\bm{\phi}})$. The resulting empirical best predictor (EBP) of the poverty measure $F_{\alpha i}$ is:
@@
Note that the EBP method from Molina and Rao (2010) requires linking survey and census households to identify sample indices, which may not be feasible in real-world applications. Often, the survey sample is not a subset of the census. To address this limitation, we follow the approach proposed by Rodas, Molina and Nguyen (2021) and apply Monte Carlo simulations to all population units rather than restricting it to non-sampled units. Thus, the EBP can be expressed as:
@@
For consistency, and without causing confusion, we use (9) as the EBP for the poverty measures throughout this paper. When the sampling fraction $n_{i}/N_{i}$ is negligible, which is the most cases in practice, the EBP (9) is practically equivalent to the original formulation proposed by Molina and Rao (2010) and further refined by Rodas, Molina and Nguyen (2021) and Molina, Corral and Nguyen (2022).
@@ 3 The LSMS and Census data
Among the observed municipalities, sample sizes vary significantly, ranging from just 6 households in municipalities such as Koder Thu and Hajmel to 600 households in Tirana. The mean sample size is 16.8, with quartiles at 8 (25th percentile), 8 (50th percentile), and 16 (75th percentile). This indicates that several municipalities have very small sample sizes, which hinders the reliability of direct estimates.
@@
where $w_{ij}$ represents the survey weight for individual $j$ in municipality $i$, and $\hat{N}_{i}=\sum_{j\in s_{i}}w_{ij}$ is a design-unbiased estimator of the population size $N_{i}$. The weighted direct estimates and their variances $\widehat{\text{Var}}(\hat{F}_{\alpha i}^{w})$ are computed using the direct function from the emdi R package.
Figure 1 illustrates the percent coefficient of variation (CV) of direct estimates for HCR and PG, emphasizing the increasing uncertainty in estimates for municipalities with smaller sample sizes. Statistical agencies set thresholds for the reliability of estimates based on their coefficient of variation (CV). For instance, Statistics Canada considers an estimate to be publishable if its CV is below $16.6\%$, whereas estimates with a CV exceeding $33.3\%$ are deemed unpublishable (Statistics Canada, 2015). Similarly, the UK Office for National Statistics (ONS) sets a threshold of 20% for publication suitability (ONS, 2020). If we set the reliability threshold at 20%, direct survey estimates of the 2002 LSMS dataset would be publishable for only 9 out of the 213 municipalities for HCR and just 2 for PG. This finding underscores the limitations of direct estimation methods in small domains, where high variability often renders estimates unreliable.
@@
Figure 1: Histograms of Percent coefficient of variation of direct estimates for HCR and PG for all sampled municipalities using 2002 LSMS data.
@@
The application of SAE methods based on the EBP significantly increases the number of municipalities for which reliable estimates can be obtained. This approach requires fitting an appropriate model to the survey data. Estimated model parameters for fixed and random effects are then combined with known population information for each municipality to predict its levels of HCR and PG. The selection of covariates is based on prior poverty assessment studies in Albania, ensuring consistency with previous methodologies. We employ the same household-level covariates as Betti, Neri and Bellini (2003) and Tzavidis et al. (2008), utilizing information available from both the survey and Census. Specifically, the following variables are considered: household size (a strong predictor of poverty), dwelling facilities (e.g., presence of TV, satellite dish, refrigerator, air conditioning, personal computer), ownership variables (e.g., homeownership, land ownership, car ownership).
@@
The descriptive statistics highlight a strong consistency between LSMS and Census data across most covariates, reinforcing the validity of using both sources for small area estimation. Household size exhibits nearly identical distributions, with the same median and very similar mean values, suggesting that the sample adequately represents population characteristics. Similarly, ownership variables such as homeownership and land ownership display minimal differences between the two sources, indicating a high degree of reliability in their measurement. Although minor discrepancies arise in asset ownership, such as car ownership and household appliances, these differences are within an expected range and may be attributed to reporting variations between the survey and the Census. Notably, the proportion of households with a television is slightly higher in LSMS than in the Census ($0.95$ vs. $0.90$), while similar patterns are observed for refrigerator and satellite dish ownership. These variations could reflect differences in reporting behavior or coverage between the two data sources. Overall, the alignment of key covariates between the two datasets provides confidence in their joint use for poverty estimation and underscores the representativeness of the survey data in capturing relevant socioeconomic conditions.
4 Nested error linear regression model with high dimensional parameters

### v2

# Empirical Best Prediction of Poverty Indicators via Nested Error Regression with High-Dimensional Parameters

@@ Abstract
The Nested Error Regression Model with High-Dimensional Parameters (NERHDP) is extended to address challenges in small area poverty estimation. A robust and flexible framework is proposed to derive empirical best predictors (EBPs) of small area poverty indicators while accommodating heterogeneity in regression coefficients and sampling variances across areas. To mitigate the computational limitations of the existing algorithm, an efficient estimation procedure is introduced, substantially reducing computation time and enhancing scalability for large datasets. A novel approach for generating area-specific poverty estimates in out-of-sample areas is also developed, improving the reliability of synthetic estimates. Uncertainty is quantified through a parametric bootstrap method specifically tailored to the extended model. Under heterogeneous data-generating scenarios, the proposed method yields lower relative bias and relative root mean squared prediction error than existing approaches. The methodology is further illustrated using data from the 2002 Albania Living Standards Measurement Survey, combined with auxiliary information from the 2001 census, to estimate poverty indicators for 374 municipalities.
Keywords: Poverty mapping, small area estimation, data integration, robust inference
@@ 1 Introduction
The United Nations emphasizes the priority of eradicating all forms of poverty by making it the first of its 17 Sustainable Development Goals. Measuring poverty levels across regions is essential for designing effective poverty reduction strategies and ensuring equitable allocation of resources. Accurate information about people’s living conditions at a regional level is a basic instrument for targeting policies and programs to reduce poverty. However, progress toward this goal is often slowed by a consistent lack of data on important economic indicators, especially for small population groups and specific geographic areas in developing countries. One major reason for the shortage of disaggregated poverty data is that large scale sample surveys are typically designed to produce reliable estimates of various characteristics of interest for large geographic areas or large population subgroups, rather than for smaller, more granular domains (Chatterjee et al., 2008).
@@
Many measures of poverty and inequality are nonlinear functions of the small area distribution of the economic variable underpinning poverty, usually income or a surrogate such as expenditure. In this paper, we focus on estimating a class of poverty measures known as the Foster–Greer–Thorbecke (FGT) poverty measures, introduced by Foster et al. (1984). Consider a finite population of size $N$ partitioned into $m$ small areas of size $N_{1},\cdots,N_{m}$. Let $E_{ij}$ be a suitable quantitative measure of welfare for individual $j$ in small area $i$, such as income or expenditure, and let $z$ be a fixed poverty line. This poverty line $z$ serves as the threshold below which an individual is considered to be living in poverty. The FGT family of poverty measures for each small area $i$ is then defined as the area-specific mean
@@
The small area estimation (SAE) literature for poverty measures was revolutionized by the pioneering work of Elbers et al. (2003), which introduced a unit-level approach (ELL) for poverty mapping. This method has become widely adopted for producing fine-grained estimates of FGT measures. The ELL approach assumes a two-level model with a common regression relationship between the response and auxiliary variables while allowing for heteroskedastic unit-level errors. Following the framework proposed by Elbers et al. (2003), the unit-level variance components are estimated through a parametric specification of heteroscedasticity as a function of covariates, which may differ from the covariates usued in the mean regression model. In later literature, this variance modeling approach has been referred to as an “alpha model” (see, e.g., Nguyen et al., 2017, Das and Chambers, 2024). Simulated census data are then generated to estimate poverty measures. It is important to note that the ELL approach is a synthetic method that does not rely on a conditional distribution. For further details, we refer readers to their original paper.
@@
Another popular approach for poverty estimation introduced by Molina and Rao (2010) is the empirical best prediction (EBP) method through Monte Carlo simulations. This method assumes that the transformed population welfare variables is characterized by a homoskedastic nested error regression (NER) model, as proposed by Battese et al. (1988). Specifically, it assumes that the regression coefficients and variance components are the same across all small areas, with normally distributed errors. Model parameters are estimated using the maximum likelihood (ML) method or the restricted maximum likelihood (REML) method, and EBPs of poverty indicators are then approximated through random draws from the conditional distribution of the out-of-sample units, given the sample data. Molina et al. (2014) introduced a Bayesian version of this approach, which is particularly useful for quantifying the uncertainty of the estimates. More recently, Morales et al. (2021) (Chapter 10) showed that the conditional expectation can be calculated explicitly and presented closed-form EBPs of the headcount ratio and poverty gap under the NER model.
@@
The effectiveness of such EBP method tends to decline when the number of small areas to be combined becomes large. This limitation arises because the assumption of identical regression coefficients and/or variance components across all small areas, inherent in these models, may not hold universally, leading to potential issues with model misspecification. To address this, random area-specific regression coefficient models (Prasad and Rao, 1990; Hobza and Morales, 2013; Rao and Molina, 2015) and random area-specific sampling variance models (Otto and Bell, 1995; Arora et al., 1997; Kubokawa et al., 2016; Sugasawa and Kubokawa, 2017) have been suggested in the literature. Such modeling, though useful in some applications, needs more nontrivial assumptions on the joint distribution for a large number of random effects. On the other hand, fixed effects assumptions on the area specific regression coefficients and sampling variances generally lead to unstable estimates of these fixed effects due to small area specific sample sizes (Jiang and Nguyen, 2012).
@@
Lahiri and Salvati (2023) introduced a novel approach in the SAE literature, addressing several challenges. Specifically, this approach assumes fixed effects for both area-specific regression coefficients and sampling variances. But they estimate these coefficients using area-specific (with respect to a tuning parameter) estimating equations applied to data from all areas, and then use appropriately constructed residuals for estimation of variance components. When the tuning parameters are known, they have shown that the model parameters can be consistently estimated. This approach, termed as the Nested Error Regression Model with High-Dimensional Parameters (NERHDP), not only offers a more flexible SAE modeling framework but also provides a reliable solution to the challenges of high-dimensional parameter estimation in small area analysis. However, their study focused solely on EBPs for small area means, which are linear functions of the population vector of outcome variable.
@@
In this paper, we extend the NERHDP model to obtain EBPs of poverty indicators, as particular complex and nonlinear FGT measures. Unlike the approach in Molina and Rao (2010), our proposed EBPs account for heterogeneity across small areas in both regression coefficients and sampling variances. Such heterogeneity is common in real-world applications due to variations in socio-economic conditions, data quality, and sampling designs among areas. By extending NERHDP to accommodate FGT poverty measures, we enable the generation of more precise and reliable estimates that are sensitive to area-specific characteristics. With suitable data, this innovation not only enhances the flexibility and applicability of the modeling but also provides a robust methodological framework for addressing the complexities inherent in small area poverty estimation.
@@
Additionally, we address the computational challenges associated with the algorithm proposed by Lahiri and Salvati (2023), which involves iterative convergence and can be computationally intensive. We introduce a more efficient algorithm that significantly reduces computation time relative to the approach in Lahiri and Salvati (2023), producing estimates within seconds. This improvement ensures the proposed model is not only methodologically robust but also practical for large-scale applications.
@@
In what follows, we focus on two widely used indicators of poverty: the Headcount Ratio (HCR), defined as the proportion of individuals below the poverty line, and the Poverty Gap (PG), which measures the average shortfall of the poor from the poverty line. We first conduct model-based simulation studies to evaluate the performance of our proposed EBP method under the NERHDP model across different data-generating processes. The simulation results show that given suitable data, our method outperforms existing predictors in terms of relative bias and relative root mean squared prediction error, when there is heterogeneity in regression coefficients and/or sampling variances. When regression coefficients are constant across areas, the performance of our proposed method is comparable to that of the EBP introduced in Molina and Rao (2010).
In addition, we demonstrate the benefits of implementing the proposed approach on data from the 2002 Living Standards Measurement Survey (LSMS) in Albania to estimate HCRs and PGs for the 374 municipalities, using the limited auxiliary information available from the 2001 Census. Of these municipalities, $161$ are out-of-sample, meaning they were not included in the original probability sample used for estimation. Consequently, their characteristics were not directly observed in the survey, and any estimates for these areas rely entirely on model-based predictions.
The structure of the paper is as follows. Section 2 provides a brief review of the EBP methodology for FGT measures. Section 3 introduces the 2002 Living Standards Measurement Survey (LSMS) in Albania, describes the estimation problem, and outlines the available auxiliary information. Section 4 introduces the best prediction approach under a nested error regression model with high-dimensional parameters, along with the associated new parameter estimation algorithms. In Section 5, we describe the parametric bootstrap method used to estimate measures of uncertainty. Section 6 presents the results of model-based simulation studies, comparing the performance of the proposed method with existing approaches in terms of bias and variability. In Section 7, we apply the new methodology to LSMS data to estimate HCR and PG at the municipality level. Finally, Section 8 concludes the paper by summarizing the key findings and suggesting directions for future research.
@@ 2 A review of Empirical Best Prediction for FGT measures
Like in Molina and Rao (2010), we assume that a random sample of size $n<N$ is drawn from the finite population using a specified sampling design. Let $\Omega$, $s\subset\Omega$, and $r=\Omega-s$ denote the set of indices for all population units, the set of sampled units, and the set of indices for the unsampled units with size NnN-n, respectively. For area $i$, we have the following associated notations: $\Omega_{i}$, $s_{i}$, and $n_{i}$. Note that if area $i$ is not sampled, then $n_{i}=0$.
@@
where the expectation is taken with respect to the conditional distribution of $\mathbf{y_{r}}$ given the observed sample data $\mathbf{y}_{s}$ under an assumed model on $\mathbf{y}$.
@@
$\hat{F}^{B}_{\alpha ij}=\text{E}_{\mathbf{y}_{r}}[h_{\alpha}(Y_{ij})|\mathbf{y}_{s}]=\int h_{\alpha}(t)f_{Y_{ij}}(t|\mathbf{y}_{s})dt,$ (5)
@@
Due to the complexity of $h_{\alpha}(t)$, it may not be always feasible to explicitly calculate the expectation in (5). Assuming that $\mathbf{y}=(\mathbf{y}_{s}^{\prime},\mathbf{y}_{r}^{\prime})^{\prime}$ follows a multivariate normal distribution with mean vector $\bm{\mu}=(\bm{\mu}_{s}^{\prime},\bm{\mu}_{r}^{\prime})^{\prime}$ and a covariance matrix partitioned as:
@@
$\mathbf{y}_{r}|\mathbf{y}_{s}\sim\mathcal{N}(\bm{\mu}_{r|s},\mathbf{V}_{r|s}),$ (6)
@@
It is important to note that the normality here is adopted as a working modeling assumption to obtain a tractable form for the conditional expectation that defines the empirical best predictor. In applied settings, this assumption may only hold approximately. However, it is standard in EBP methodology and provides a convenient parametric basis for prediction and uncertainty assessment (see, e.g., Molina and Rao, 2010; Rao and Molina, 2015).
@@
[…]
@@
Note that the EBP method from Molina and Rao (2010) requires linking survey and census households to identify sample indices, which may not be feasible in real-world applications. Often, the survey sample is not a subset of the census. To address this limitation, we follow the approach proposed by Rodas et al. (2021) and apply Monte Carlo simulations to all population units rather than restricting it to non-sampled units. Thus, the EBP can be expressed as:
@@
For consistency, and without causing confusion, we use (9) as the EBP for the poverty measures throughout this paper. When the sampling fraction $n_{i}/N_{i}$ is negligible, which is the most cases in practice, the EBP (9) is practically equivalent to the original formulation proposed by Molina and Rao (2010) and further refined by Rodas et al. (2021) and Molina et al. (2022).
@@ 3 The LSMS and Census data
Among the observed municipalities, sample sizes vary significantly, ranging from just 6 households in municipalities such as Koder Thu and Hajmel to 600 households in Tirana. The mean sample size is 16.8, with quartiles at 8 (25th percentile), 8 (50th percentile), and 16 (75th percentile). This indicates that several municipalities have very small sample sizes, which hinders the reliability of direct estimates and the stable estimation of random-effect variance components in model-based approaches.
@@
where $w_{ij}$ represents the survey weight for individual $j$ in municipality $i$, and $\hat{N}_{i}=\sum_{j\in s_{i}}w_{ij}$ is a design-unbiased estimator of the population size $N_{i}$. The weighted direct estimates and their variances, $\widehat{\text{Var}}_{D}(\hat{F}_{\alpha i}^{w})$, are computed using the direct function from the emdi R package, which employs a nonparametric bootstrap procedure to estimate the design-based variance of the direct estimator $\hat{F}_{\alpha i}^{w}$. This approach is consistent with the general bootstrap methods for complex surveys discussed in the literature (see, e.g., Rao and Molina (2015), Chapter 5).
Figure 1 illustrates the average percent estimated coefficient of variation (CV %) of direct estimates for HCR and PG, emphasizing the increasing uncertainty in estimates for municipalities with smaller sample sizes.
@@
Figure 1: Bar chart of average percent estimated coefficient of variation of direct estimates for HCR and PG versus within-area sample size for all sampled municipalities.
Statistical agencies set thresholds for the reliability of estimates based on their CVs. For instance, Statistics Canada considers an estimate to be publishable if its CV is below $16.6\%$, whereas estimates with a CV exceeding $33.3\%$ are deemed unpublishable (Statistics Canada, 2015). Similarly, the UK Office for National Statistics (ONS) sets a threshold of 20% for publication suitability (ONS, 2020). If we set the reliability threshold at 20%, direct survey estimates of the 2002 LSMS dataset would be publishable for only 9 out of the 213 municipalities for HCR and just 2 for PG. Further details are provided in Figure 2, which displays the histograms of the estimated percent coefficients of variation for the direct estimates of HCR and PG across all sampled municipalities in the 2002 LSMS data.
This result highlights the limitations of direct estimation methods in small domains, where large sampling variability often renders estimates unreliable. It is important to note that the coefficient of variation is only appropriate for strictly positive outcome variables, such as the poverty indicators considered in this paper (HCR and PG), which are bounded between 0 and 1. By contrast, variables such as household income or expenditure, which may take non-positive values, require alternative measures of reliability. In such cases, statistical agencies and the SAE literature recommend the use of the standard error (SE) or the relative root mean squared prediction error (RRMSPE) as more suitable indicators of precision (see, e.g., Rao and Molina (2015)). Accordingly, in our subsequent simulation studies, we report RRMSPE alongside relative bias to evaluate the performance of predictors.
Figure 2: Histograms of percent estimated coefficient of variation of direct estimates for HCR and PG for all sampled municipalities using 2002 LSMS data.
@@
The application of SAE methods based on the EBP significantly increases the number of municipalities for which reliable estimates can be obtained. This approach requires fitting an appropriate model to the survey data. Estimated model parameters for fixed and random effects are then combined with known population information for each municipality to predict its levels of HCR and PG. The selection of covariates is based on prior poverty assessment studies in Albania, ensuring consistency with previous methodologies. We employ the same household-level covariates as Betti et al. (2003) and Tzavidis et al. (2008), utilizing information available from both the survey and Census. Specifically, the following variables are considered: household size (a strong predictor of poverty), dwelling facilities (e.g., presence of TV, satellite dish, refrigerator, air conditioning, personal computer), ownership variables (e.g., homeownership, land ownership, car ownership).
@@
The descriptive statistics highlight a strong consistency between LSMS and Census data across most covariates, reinforcing the validity of using both sources for small area estimation. Household size exhibits nearly identical distributions, with the same median and very similar mean values, suggesting that the sample adequately represents population characteristics. Similarly, ownership variables such as homeownership and land ownership display minimal differences between the two sources, indicating a high degree of reliability in their measurement. Although minor discrepancies arise in asset ownership, such as car ownership and household appliances, these differences are within an expected range and may be attributed to reporting variations between the survey and the Census. Notably, the proportion of households with a television is slightly higher in LSMS than in the Census ($0.95$ vs. $0.90$), while similar patterns are observed for refrigerator and satellite dish ownership. These variations could reflect differences in reporting behavior or coverage between the two data sources. Overall, the alignment of key covariates between the two datasets provides confidence in their joint use for poverty estimation and underscores the representativeness of the survey data in capturing relevant socioeconomic conditions. It is worth noting that auxiliary variables expressed as proportions that are close to zero or one may have limited predictive value due to their constrained variability. In practice, the contribution of such variables to small area models should be assessed in light of their dispersion across areas.
4 Nested error regression model with high dimensional parameters